# Running in circles: is practical application feasible for data fission and data thinning in post-clustering differential analysis?

### v1

@@ 2.3 Practical solutions
2.3.1 Individual fission (or thining) for mixture models
In Gaussian mixture models, parameters are typically component-specific, meaning that they are assumed to be shared across all individuals within each component. As explained above and in Figure 1, this assumption poses a circular challenge as it requires knowledge of the true data structure to be able to accurately estimate the component-specific covariance matrices that are then needed to perform data fission or data thinning. To address this limitation, we propose an alternative approach in which each observation is modeled as a realization of its own Gaussian distribution, i.e. $p_{i}(\boldsymbol{x}_{i})=f\left(\boldsymbol{x_{i}}|\boldsymbol{\mu}_{i},\boldsymbol{\Sigma}_{i}\right)$. Thus, observations are no longer assumed to be identically distributed). Consequently, the covariance matrix $\boldsymbol{\Sigma}_{i}$ is no longer specific to the components but instead to the individual observations. Despite this individual-level definition, two individuals drawn from the same component are expected to have very close variance parameters. By bypassing the components in the definition of $\boldsymbol{\Sigma}_{i}$, this modeling strategy theoretically encompasses both the global null ($K=1$) and the mixture ($K\geq 1$) settings, and opens up individual-level data fission and data thinning. As highlighted above, variance estimation remains crucial for the practical feasibility of these methods. This new modeling assumes individual variances that still need to be known (ideally), or precisely estimated in real-life settings.
@@ 2.3.2 Non-parametric local variance estimation
The catch of the variance computation in (2) lies in determining each individual weight $w_{ij}$. Since it is essential for $w_{ij}$ to appropriately capture the proximity between observations, we opted for a kernel-based definition: $w_{ij}=K\left(x_{i}-x_{j}\right)$, where $K(u)=\frac{1}{\sqrt{2\pi}h}e^{-\frac{u^{2}}{2h^{2}}}$ represents the uniform kernel providing the proximity measurement between $x_{i}$ and $x_{j}$. This ensures that individual weights reflect the local relationships within the data. This kernel choice focuses on nearby points, emphasizing their influence on the weighted variance estimate. The parameter $h$ in the definition of $K$ serves as the bandwidth parameter, controlling the width of the kernel and, consequently, the neighborhood around each observation $x_{i}$ that contributes to its weighted variance estimate. A smaller $h$ results in more localized estimations, emphasizing nearby points, but can lead to underestimate the variances. Conversely, a larger $h$ includes a broader range of observations, and an excessively large value might consider almost all observations in the variance estimation, yielding an estimator close to the full sample one. Therefore, an optimal choice for $h$ would involve considering only the observations in the same components of the underlying mixture. However, achieving this ideal scenario is impractical as it is equivalent to knowing the true components of the mixture.
Bandwidth calibration is a crucial step in any kernel method (Heidenreich et al., 2013). We propose to use individual-specific bandwidth $h_{i}=h(x_{i})$ to reduce the bias in kernel density estimation. To accomplish this, we first estimate the changepoint, $i^{*}$ in the spread of distances from observation $x_{i}$ to all other observations. This changepoint delineates the change in components of the mixture: observations with distances preceding this changepoint are deemed part of the same component as $x_{i}$, whereas those following it are considered part of different components. Subsequently, we define $h_{i}=|x_{i}-x_{i^{*}}|$ as the distance between $x_{i}$ and the observation $x_{i^{*}}$ that is furthest from $x_{i}$ before the break in the mixture. Thus, $h_{i}$ is determined in such a way that the individual bandwidth accommodates only the observations in the same component as $x_{i}$, ensuring its size is sufficient to encompass all observations of the component (Chacón & Duong, 2020).
@@ 3.1 Quantification of Type I error rate as a function of the bias in the variance estimation
Independence between $\boldsymbol{X}^{(1)}$ and $\boldsymbol{X}^{(2)}$ is guaranteed solely when the true variance is used for decomposition. Substituting the true variance $\boldsymbol{\Sigma}_{i}$ with an estimate $\widehat{\boldsymbol{\Sigma}}_{i}$ introduces correlations between these new random variables. Consider a Gaussian random variable $\boldsymbol{X}\sim\mathcal{N}\left(\boldsymbol{\mu}_{p},\boldsymbol{\Sigma}_{p\times p}\right)$ that we aim to decompose using the data fission or data thinning processes outlined in Table 1, but with an estimated covariance matrix, $\widehat{\boldsymbol{\Sigma}}$, in place of $\boldsymbol{\Sigma}$. Table 2 provides the covariance values between $\boldsymbol{X}^{(1)}$ and $\boldsymbol{X}^{(2)}$ as a function of the bias of $\widehat{\boldsymbol{\Sigma}}$ (for proofs, refer to Appendix B for data fission and to Neufeld, Dharamshi, Gao & Witten (2023) for data thinning). If $\boldsymbol{X}^{(1)}$ and $\boldsymbol{X}^{(2)}$ are not independent (as with a biased estimator of $\boldsymbol{\Sigma}$ that induces significant covariance), a split of a single component into two estimated clusters in $\boldsymbol{X}^{(1)}$ is easily transferred to $\boldsymbol{X}^{(2)}$ and results in false positives during the inference step (even if the latter is carried out on $\boldsymbol{X}^{(2)}$).
Data Fission Data Thinning
$\operatorname{Cov}\left(\boldsymbol{X}^{(1)},\boldsymbol{X}^{(2)}\right)$ $\boldsymbol{\Sigma}-\widehat{\boldsymbol{\Sigma}}$ $\tau(1-\tau)\left(\boldsymbol{\Sigma}-\widehat{\boldsymbol{\Sigma}}\right)$
Table 2: Bias in the estimation of $\boldsymbol{\Sigma}$ induces correlations between $\boldsymbol{X}^{(1)}$ and $\boldsymbol{X}^{(2)}$
@@
To further describe the repercussions of variance estimation in the context of data fission, we derived an analytical expression for the Type I error rate of the $t$-test as a function of the bias in estimating this parameter. Let $X\sim\mathcal{N}(\mu,\sigma^{2})$ be a random variable. Given that $X$ is not generated from a mixture of distributions, it becomes evident that distinct clusters are not inherent in the variable. Our aim is to perform data fission in $X$, using $X^{(1)}$ for $k-$means clustering (with $K=2$) and $X^{(2)}$ for differential testing between the two inferred clusters. Notably, we employ an estimate $\widehat{\sigma^{2}}$ of $\sigma^{2}$ in the fission process. Given the generation process of $X$, it is established that regardless of the clustering on $X^{(1)}$, there should be no mean difference between the estimated clusters on $X^{(2)}$ as long as independence is achieved. Let $\widehat{C}_{1}$ and $\widehat{C}_{2}$ be the two estimated clusters on $X^{(1)}$ with the same intra-cluster variance, which is a reasonable hypothesis with $k-$means clustering as explained in Appendix C. Since we are under the null hypothesis of no mean difference between the clusters, the $T$ statistic for the $t$-test between $\widehat{C}_{1}$ and $\widehat{C}_{2}$ on $X^{(2)}$ is given by:
$T_{\widehat{C}_{1},\widehat{C}_{2}}=\frac{\widehat{\delta}_{X^{(2)}}}{\sqrt{\frac{4s^{2}(X^{(2)})}{n}}}$ (3)
Here, $\widehat{\delta}_{X^{(2)}}$ represents the observed mean difference between $\widehat{C}_{1}$ and $\widehat{C}_{2}$ on $X^{(2)}$ and $s^{2}\left(X^{(2)}\right)$ is their shared intra-cluster variance. Of note, $\widehat{\delta}_{X^{(2)}}$ and $s^{2}\left(X^{(2)}\right)$ both depend on the ability to reproduce the clustering results obtained on $X^{(1)}$ into $X^{(2)}$. This is quantified by their correlation $\operatorname{Cor}\left(X^{(1)},X^{(2)}\right)$, which can be easily computed from their covariance given in Table 2. Based on this correlation, it can be demonstrated that:
$T_{\widehat{C}_{1},\widehat{C}_{2}}\stackrel{{\scriptstyle\mathcal{H}_{0}}}{{\sim}}\mathcal{N}\left(\frac{\sqrt{n}\sqrt{\frac{2}{\pi}\operatorname{Cor}\left(X^{(1)},X^{(2)}\right)^{2}}}{\sqrt{1-\frac{2}{\pi}\operatorname{Cor}\left(X^{(1)},X^{(2)}\right)^{2}}},1\right)$
@@
where $\operatorname{Cor}\left(X^{(1)},X^{(2)}\right)=\frac{\left(\sigma^{2}-\widehat{\sigma^{2}}\right)}{\sqrt{\left(\sigma^{2}+\tau^{2}\widehat{\sigma^{2}}\right)\times\left(\sigma^{2}+\frac{1}{\tau^{2}}\widehat{\sigma^{2}}\right)}}$. Appendix C gives details on the derivation of this test statistic and its distribution. The associated Type I error rate for this test is given by $1-F(q_{\alpha/2})+F(-q_{\alpha/2})$, where $F$ is the cumulative distribution function of $\mathcal{N}\left(\frac{\sqrt{n}\sqrt{\frac{2}{\pi}\operatorname{Cor}\left(X^{(1)},X^{(2)}\right)^{2}}}{\sqrt{1-\frac{2}{\pi}\operatorname{Cor}\left(X^{(1)},X^{(2)}\right)^{2}}},1\right)$ (or the corresponding non-central $\mathcal{T}(n-2)$ distribution), and $q_{\alpha/2}$ is the quantile of a standard Gaussian distribution $\mathcal{N}(0,1)$ (or of a $\mathcal{T}(n-2)$ distribution).
@@
We validated this theoretical result through numerical simulations, and conducted a detailed exploration of the influence of variance values and sample size on the resulting Type I error rate. We generated $n$ realizations of a Gaussian random variable with a mean $\mu=0$ and variances $\sigma^{2}$, with $1,\!000$ Monte Carlo repetitions. Subsequently, we used data fission with varying values of $\widehat{\sigma^{2}}$, obtaining $X^{(1)}$ for k-means clustering with $K=2$ and $X^{(2)}$ for testing mean differences between the two estimated clusters. Initially, we examine the impact of the original true variance $\sigma^{2}$ by considering $\sigma^{2}$ values of $\{0.01,1,4\}$ for a fixed sample size of $n=100$. Figure 3A, illustrates the relationship between the bias in estimating $\sigma^{2}$ and the Type I error rate, demonstrating a consistent agreement with the derived theoretical error rate. We further documented the behavior of the Type I error for a fixed $\sigma^{2}=1$ with varying sample sizes $n\in\{50,100,250,100,1000\}$ in Figure 3B, which shows the expected impact of the sample size on the Type I error rate. These findings collectively underscore the critical importance of accurately estimating the variance to achieve a well-calibrated Type I error rate with data fission.
@@ 3.2 Performances of the local variance estimator
Figure 4A shows that local variances are underestimated (manifested as a negative relative bias) until the signal vs. noise ratio $\delta/\sigma$ reaches approximately $3$ – a threshold value for separation in a Gaussian mixture model previously reported in the literature (Siffer et al., 2018; Hivert, Agniel, Thiébaut & Hejblum, 2024). Consequently, there are no clear change-points in the spread of distances for individual observations and all observations should contribute significantly to the variance calculation. However, this is not the case here (with a bandwidth $h_{i}$ that is too small), leading to underestimated local variances. As the ratio increases within the range $3.5\leq\delta/\sigma\leq 10$, component separation becomes clearer but the identification of change-points remains challenging. Change-points detection is intricate for observations in the tails in-between the components, leading to non-zero weight for observations in both components. Consequently, an overestimation of the local variances (positive relative biases) is observed together with an increase in their associated inter-quantile range. Finally, for $\delta/\sigma>10$, sufficient component separation ensures consistent local variance estimation, highlighting the critical importance of clear-cut component separation for accurate variance estimation. Thus, the methodology outlined in 2.3.2 requires well-separated observations to be able to provide accurate variance estimations. In Figure 4B, the Type I error rate at $5\%$ remains well calibrated only for values of $\delta/\sigma$ that ensures robust variance estimations (that is $\delta/\sigma>10$), emphasizing again the necessity of good component separations for reliable testing. Figure 4C demonstrates how increasing the sample size fails to improve non-parametric local variance estimations for three representative values of the ratio $\delta/\sigma\in\{0.5,5,10\}$, indicating that performance is more dependent upon component separation rather than sample size. The corresponding Type I error rates depicted in Figure 4B thus align with the observed relative bias results, underscoring the critical role of accurate estimation of local variances to ensure Type I error control.
@@ 3.3 Application to single-cell RNA-seq data analysis
Single-cell RNA-seq (scRNA-seq) data analysis pipelines often involve an initial cluster analysis followed by differential analysis to identify marker genes and annotate cell populations based on gene expression. Given their overdispersed count nature, the negative binomial distribution is favored over the Gaussian distribution to model scRNA-seq data. Unfortunately, the negative binomial distribution also raises challenges for data thinning. The overdispersion parameter $\theta$ plays a role similar to the variance parameter in the Gaussian distribution, as it is considered known. Therefore, the quality of its estimation is directly impacts $\operatorname{Cov}\left(X^{(1)},X^{(2)}\right)=\tau(1-\tau)\frac{\mu^{2}}{\theta}\left(1-\frac{\theta+1}{\widehat{\theta}+1}\right)$, where $\widehat{\theta}$ is an estimated of $\theta$ (Neufeld, Dharamshi, Gao & Witten, 2023). Also, for negative binomial mixtures, data thinning decomposition is once again feasible only at the component level. As the overdispersion parameter is component-specific (Li et al., 2018), providing an estimator that ensures independence is here also a harduous and circular task, given that the components themselves are again unknown and thus require estimation through data thinning.
@@
We further illustrate the necessity of applying intra-component data thinning (with the associated intra-components overdispersion) to ensure Type I error control in the post-clustering inference setting with numerical simulations. We generated $n=100$ observations from a two-component negative binomial mixture: $0.5\text{NegBin}(\mu_{1},\theta_{1})+0.5\text{NegBin}(\mu_{2},\theta_{2})$ with component parameters $(\mu_{1},\theta_{1})=(5,5)$ and $(\mu_{2},\theta_{2})=(60,40)$. We conducted similar post-clustering inference as in Figure 2. In Figure 5A, applying the $k-$means with $K=3$ clusters reveals that the first mixture component is erroneously split into 2 clusters (labeled $C_{1}$ and $C_{3}$). We then assessed the Type I error rate associated with the Wilcoxon test between these incorrect clusters when applying data thinning with various overdispersion estimators. First, we applied intra-component data thinning using oracle estimates $\widetilde{\theta}_{k}$, $k=1,2$, representing true intra-component overdispersion parameters (unfeasible in real-life application where the true cluster structure is unknown). We compared those results with two alternatives that are feasible in practice: i) applying intra-cluster data thinning based on the $k-$means results from Figure 5A with their associated $\widehat{\theta}_{\hat{k}}$, $\hat{k}=1,2,3$, and global data thinning with its associated $\widehat{\theta}$. Of note, all the overdispersion estimations were performed using Maximum Likelihood. Figure 5B presents the QQ-plot against the Uniform distribution of the associated Wilcoxon p-values across $1,\!000$ simulations. Similarly to the Gaussian setting, achieving uniformly distributed p-values is only possible through intra-component data thinning using their oracle overdispersion estimates $\widetilde{\theta}_{k}$. All other data thinning approaches are performed with biased estimators of the overdispersion compromising the independence between $X^{(1)}$ and $X^{(2)}$, and therefore leading to failure in the control of the post-clustering Type I error rate.
@@
Extending our investigation from simulated scenarios to real-life applications, we used a single-cell RNA-seq dataset from the Tabula Sapiens Consortium (Consortium* et al., 2022) to delve into the practical challenges of estimating overdispersion. Our analysis focused on five distinct cell populations: $2,\!560$ neutrophils, $105$ macrophages, $386$ monocytes, $454$ granulocytes, and $833$ CD4 T cells – all collected from a single donor. In this controlled setting, where cell types were known, we succeeded in estimating the overdispersion of $8,\!333$ genes for each cell type using the variance stabilizing transformation implemented in the sctransform package (Choudhary & Satija, 2022). Figure 5C illustrates the comparison of gene overdispersion when estimated solely in neutrophils versus the overdispersion estimated for the same genes in the other four cellular populations. Root Mean Squared Error (RMSE) values were computed to quantify the agreement between estimations. Our findings reveal that overdispersion is specific to each cell population, as evidenced by a notable deviation from the diagonal and relatively high RMSE values. This underscores the challenge of accurately estimating this parameter without prior knowledge of the true mixture underlying the data. Combined with our simulation studies, these results demonstrate how Type I error can easily be inflated in real-life applications of data thining for scRNA-seq data analysis, particularly due to the difficulty in providing an unbiased estimate of gene overdispersion.

### v2

@@ 2.3 Practical solutions
To address the circular situation of data fission and thinning in Gaussian mixture models, we propose modeling each observation as a realization of its own Gaussian distribution. This approach bypasses the need for prior knowledge of the true data structure, allowing for individual-level data fission and thinning. Nonetheless, the accurate estimation of individual variance parameters remains a critical challenge for practical feasibility.
2.3.1 Individual fission (or thinning) for mixture models
In Gaussian mixture models, parameters are typically component-specific, meaning that they are assumed to be shared across all individuals within each component. As explained above and in Figure 1, this assumption poses a circular challenge as it requires knowledge of the true data structure to be able to accurately estimate the component-specific covariance matrices that are then needed to perform data fission or data thinning. To address this limitation, we propose an alternative approach where each observation is modeled as a realization of its own Gaussian distribution, i.e. $p_{i}(\boldsymbol{x}_{i})=f\left(\boldsymbol{x_{i}}|\boldsymbol{\mu}_{i},\boldsymbol{\Sigma}_{i}\right)$. Thus, observations are no longer assumed to be identically distributed. Consequently, the covariance matrix $\boldsymbol{\Sigma}_{i}$ is no longer specific to the components but instead to the individual observations. Despite this individual-level definition, two individuals drawn from the same (unknown) component are expected to have very close variance parameters. By bypassing the components in the definition of $\boldsymbol{\Sigma}_{i}$, this modeling strategy theoretically encompasses both the global null ($K=1$) and the mixture ($K\geq 1$) settings, and opens up individual-level data fission and data thinning. As highlighted above, variance estimation remains crucial for the practical feasibility of these methods. This new modeling assumes individual variances that still need to be known (ideally), or precisely estimated in real-life settings.
@@ 2.3.2 Non-parametric local variance estimation
The catch of the variance computation in (2) lies in determining each individual weight $w_{ij}$. Since it is essential for $w_{ij}$ to appropriately capture the proximity between observations, we opted for a kernel-based definition: $w_{ij}=K\left(x_{i}-x_{j}\right)$, where $K(u)=\frac{1}{\sqrt{2\pi}h}e^{-\frac{u^{2}}{2h^{2}}}$ represents the Gaussian kernel providing the proximity measurement between $x_{i}$ and $x_{j}$. This ensures that individual weights reflect the local relationships within the data. This kernel choice focuses on nearby points, emphasizing their influence on the weighted variance estimate. The parameter $h$ in the definition of $K$ serves as the bandwidth parameter, controlling the width of the kernel and, consequently, the neighborhood around each observation $x_{i}$ that contributes to its weighted variance estimate. A smaller $h$ results in more localized estimations, emphasizing nearby points, but can underestimate the variances. Conversely, a larger $h$ includes a broader range of observations, and an excessively large value might consider almost all observations in the variance estimation, yielding an estimator close to the full sample one $\mathbf{\widehat{\Sigma}}$. Therefore, an optimal choice for $h$ would involve considering only the observations in the same components of the underlying mixture. However, achieving this ideal scenario is impractical as it is equivalent to knowing the true components of the mixture.
Bandwidth calibration is a crucial step in any kernel method (Heidenreich et al.,, 2013). We propose to use individual-specific bandwidth $h_{i}=h(x_{i})$ to reduce the bias in kernel density estimation. To accomplish this, we first estimate the changepoint, $i^{*}$ in the spread of distances from observation $x_{i}$ to all other observations. This changepoint delineates the change in components of the mixture: observations with distances preceding this changepoint are deemed part of the same component as $x_{i}$, whereas those following it are considered part of different components. Subsequently, we define $h_{i}=|x_{i}-x_{i^{*}}|$ as the distance between $x_{i}$ and the observation $x_{i^{*}}$ that is furthest from $x_{i}$ before the break in the mixture. Thus, $h_{i}$ is determined in such a way that the individual bandwidth accommodates only the observations in the same component as $x_{i}$, ensuring its size is sufficient to encompass all observations of the component (Chacón and Duong,, 2020).
@@ 3.1 Quantification of Type I error rate as a function of the bias in the variance estimation
Independence between $\boldsymbol{X}^{(1)}$ and $\boldsymbol{X}^{(2)}$ is guaranteed solely when the true variance is used for decomposition. Substituting the true variance $\boldsymbol{\Sigma}_{i}$ with an estimate $\widehat{\boldsymbol{\Sigma}}_{i}$ introduces correlations between these new random variables as demonstrated by Proposition 1 for data fission (a proof of this proposition could be found in Section $2$ of the Supplementary Materials) and by Proposition $10$ of Neufeld et al., 2023a .
Proposition 1
Let $\boldsymbol{X}\sim\mathcal{N}\left(\boldsymbol{\mu}_{p},\boldsymbol{\Sigma}_{p\times p}\right)$ be a Gaussian random variable. Suppose we apply data fission on $\boldsymbol{X}$ as described in Table 1 but using $\boldsymbol{Z}\sim\mathcal{N}\left(\boldsymbol{0},\widehat{\boldsymbol{\Sigma}}\right)$, where $\widehat{\boldsymbol{\Sigma}}$ is an estimate of the $\boldsymbol{\Sigma}$. Then, $\operatorname{Cov}\left(\boldsymbol{X}^{(1)},\boldsymbol{X}^{(2)}\right)=\boldsymbol{\Sigma}-\widehat{\boldsymbol{\Sigma}}$.
Proposition 1 indicates that if $\boldsymbol{X}^{(1)}$ and $\boldsymbol{X}^{(2)}$ are not independent, as with a biased estimator of $\boldsymbol{\Sigma}$ that induces significant covariance, a split of a single component into two estimated clusters in $\boldsymbol{X}^{(1)}$ is easily transferred to $\boldsymbol{X}^{(2)}$ and results in false positives during the inference step (even if the latter is carried out on $\boldsymbol{X}^{(2)}$).
@@
To further describe the repercussions of variance estimation in the context of data fission, we derived an analytical expression for the Type I error rate of the $t$-test as a function of the bias in estimating this parameter. Let $X_{1},\dots,X_{n}$ be $n$ independent and identically distributed random variables such that, for all $i=1,\dots,n$, $X_{i}\overset{i.i.d}{\sim}\mathcal{N}\left(\mu,\sigma^{2}\right)$. Here, $n$ represents the sample size. Recall that the sample $X_{1},\dots,X_{n}$ is normally distributed: it contains no real clusters, and therefore, no actual difference in means exists between subgroups of observations that is not a consequence of the clustering. Let $Z_{1},\dots,Z_{n}\overset{i.i.d}{\sim}\mathcal{N}\left(0,b^{2}\right)$ and $\tau\in,]0,+\infty[$. Our aim is to perform data fission of each $X_{i}$, using the $X^{(1)}_{i}$ for $k-$means clustering (with $K=2$) and the $X^{(2)}_{i}$ for differential testing between the two inferred clusters. Here, $b^{2}$ represents any value used as a plug-in for the variance of the $X_{i}$, and in particular, $b^{2}$ may be an estimate obtained for $\sigma^{2}$. Given the generation process of the $X_{i}$, it is established that regardless of the clustering on the $X^{(1)}_{i}$, there should be no mean difference between the estimated clusters on $X^{(2)}_{i}$ as long as independence is achieved. Let $C_{1}$ and $C_{2}$ be the two estimated clusters on the $X^{(1)}_{i}$ with the same intra-cluster variance, which is a reasonable hypothesis with $k-$means clustering as explained in Section $3$ of the Supplementary Materials. Since we are under the null hypothesis of no mean difference between the clusters, the $T$ statistic for the $t$-test between $C_{1}$ and $C_{2}$ using the $X^{(2)}_{i}$ is given by:
$\frac{\overline{X^{(2)}_{C_{1}}}-\overline{X^{(2)}_{C_{2}}}}{\sqrt{\frac{4s^{2}\left(X^{(2)}\right)}{n}}}\qquad\mbox{where}\qquad\overline{X^{(2)}_{C_{j}}}=\frac{1}{\left|C_{k}\right|}\sum\limits_{i\in C_{k}}X_{i}^{(2)}$ (3)
Here, $s^{2}\left(X^{(2)}\right)$ is their shared intra-cluster variance computed using the $X^{(2)}_{i}$. It can be demonstrated that:
$T\overset{\mathcal{L}}{\sim}\mathcal{N}\left(\frac{\rho\sqrt{n}}{\sqrt{\frac{\pi}{2}-\rho^{2}}},1\right)$
@@
where $\rho=\operatorname{Cor}\left(X^{(1)}_{i},X^{(2)}_{i}\right)=\frac{\left(\sigma^{2}-b^{2}\right)}{\sqrt{\left(\sigma^{2}+\tau^{2}b^{2}\right)\times\left(\sigma^{2}+\frac{1}{\tau^{2}}b^{2}\right)}}$. Section $3$ of the Supplementary Materials gives details on the derivation of this test statistic and its distribution. The associated Type I error rate for this test is given by $1-F(q_{\alpha/2})+F(-q_{\alpha/2})$, where $F$ is the cumulative distribution function of $\mathcal{N}\left(\frac{\rho\sqrt{n}}{\sqrt{\frac{\pi}{2}-\rho^{2}}},1\right)$ and $q_{\alpha/2}$ is the quantile of a standard Gaussian distribution $\mathcal{N}(0,1)$.
@@
We validated this theoretical result through numerical simulations, and conducted a detailed exploration of the influence of variance values and sample size on the resulting Type I error rate. We generated $n$ realizations of a Gaussian random variable with a mean $\mu=0$ and variances $\sigma^{2}$, with $1,\!000$ Monte Carlo repetitions. Subsequently, we used data fission with varying values of $b^{2}=\widehat{\sigma^{2}}$, obtaining $X^{(1)}$ for $k-$means clustering with $K=2$ and $X^{(2)}$ for testing mean differences between the two estimated clusters. Initially, we examine the impact of the original true variance $\sigma^{2}$ by considering $\sigma^{2}$ values of $\{0.01,0.25,1,4\}$ for a fixed sample size of $n=100$. Figure 3A, illustrates the relationship between the bias in estimating $\sigma^{2}$ and the Type I error rate, demonstrating a consistent agreement with the derived theoretical error rate.
We further documented the behavior of the Type I error for a fixed $\sigma^{2}=1$ with varying sample sizes $n\in\{50,100,200,100,500,1,\!000\}$ in Figure 3B, which shows the expected impact of the sample size on the Type I error rate. These findings collectively underscore the critical importance of accurately estimating the variance to achieve a well-calibrated Type I error rate with data fission.
@@ 3.2 Performances of the local variance estimator
Figure 4A shows that local variances are underestimated (manifested as a negative relative bias) until the signal vs. noise ratio $\delta/\sigma$ reaches approximately $3$ – a threshold value for separation in a Gaussian mixture model previously reported in the literature (Siffer et al.,, 2018; Hivert et al., 2024a, ). Consequently, there are no clear change-points in the spread of distances for individual observations and all observations should contribute significantly to the variance calculation. However, this is not the case here (with a bandwidth $h_{i}$ that is too small), leading to underestimated local variances. As the ratio increases within the range $3.5\leq\delta/\sigma\leq 10$, component separation becomes clearer but the identification of change-points remains challenging. Change-points detection is intricate for observations in the tails in-between the components, leading to non-zero weight for observations in both components. Consequently, an overestimation of the local variances (positive relative biases) is observed together with an increase in their associated inter-quantile range. Finally, for $\delta/\sigma>10$, sufficient component separation ensures consistent local variance estimation, highlighting the critical importance of clear-cut component separation for accurate variance estimation. So, as long as observations are well separated, our methodology outlined in 2.3.2 is able to provide accurate variance estimates. In Figure 4B, the Type I error rate at $5\%$ remains well calibrated for those values of $\delta/\sigma$ that ensure robust variance estimation (i.e. $\delta/\sigma>10$), illustrating the reliable testing performance with good component separations. Figure 4C demonstrates how increasing the sample size does not impact the performance of the non-parametric local variance estimations for three representative values of the ratio $\delta/\sigma\in\{0.5,5,10\}$, pointing towards component separation as the main driver of testing performance. The corresponding Type I error rates depicted in Figure 4D thus align with the observed relative bias results, underscoring the critical role of accurate estimation of local variances to ensure Type I error control.
@@ 3.3 Application to single-cell RNA-seq data analysis
Single-cell RNA-seq (scRNA-seq) data analysis pipelines often involve an initial cluster analysis followed by differential analysis to identify marker genes and annotate cell populations based on gene expression. Given their overdispersed count nature, the negative binomial distribution is favored over the Gaussian distribution to model scRNA-seq data. Unfortunately, the negative binomial distribution also raises challenges for data thinning. The overdispersion parameter $\theta$ plays a role similar to the variance parameter in the Gaussian distribution, as it is considered known. Therefore, the quality of its estimation is directly impacts $\operatorname{Cov}\left(X^{(1)},X^{(2)}\right)=\tau(1-\tau)\frac{\mu^{2}}{\theta}\left(1-\frac{\theta+1}{\widehat{\theta}+1}\right)$, where $\widehat{\theta}$ is an estimated of $\theta$ (Neufeld et al., 2023a, ). Also, for negative binomial mixtures, data thinning decomposition is once again feasible only at the component level. As the overdispersion parameter is component-specific (Li et al.,, 2018), providing an estimator that ensures independence is here also a harduous and circular task, given that the components themselves are again unknown and thus require estimation through data thinning.
@@
We further illustrate the necessity of applying intra-component data thinning (with the associated intra-components overdispersion) to ensure Type I error control in the post-clustering inference setting with numerical simulations. We generated $n=100$ observations from a two-component negative binomial mixture: $0.5\text{NegBin}(\mu_{1},\theta_{1})+0.5\text{NegBin}(\mu_{2},\theta_{2})$ with component parameters $(\mu_{1},\theta_{1})=(5,5)$ and $(\mu_{2},\theta_{2})=(60,40)$. We conducted similar post-clustering inference as in Figure 2. In Figure 5A, applying the $k-$means with $K=3$ clusters reveals that the first mixture component is erroneously split into 2 clusters (labeled $C_{1}$ and $C_{3}$). We then assessed the Type I error rate associated with the Wilcoxon test between these incorrect clusters when applying data thinning with various overdispersion estimators. First, we applied intra-component data thinning using oracle estimates $\widetilde{\theta}_{k}$, $k=1,2$, representing true intra-component overdispersion parameters (infeasible in real-life application where the true cluster structure is unknown). We compared those results with two alternatives that are feasible in practice: i) applying intra-cluster data thinning based on the $k-$means results from Figure 5A with their associated $\widehat{\theta}_{\hat{k}}$, $\hat{k}=1,2,3$, and global data thinning with its associated $\widehat{\theta}$. Of note, all the overdispersion estimations were performed using Maximum Likelihood. Figure 5B presents the QQ-plot against the Uniform distribution of the associated Wilcoxon p-values across $1,\!000$ simulations. Similarly to the Gaussian setting, achieving uniformly distributed p-values is only possible through intra-component data thinning using their oracle overdispersion estimates $\widetilde{\theta}_{k}$. All other data thinning approaches are performed with biased estimators of the overdispersion compromising the independence between $X^{(1)}$ and $X^{(2)}$, and therefore leading to failure in the control of the post-clustering Type I error rate.
@@
Extending our investigation from simulated scenarios to real-life applications, we used a single-cell RNA-seq dataset from the Tabula Sapiens Consortium (Consortium* et al.,, 2022) to delve into the practical challenges of estimating overdispersion. Our analysis focused on five distinct cell populations: $2,\!560$ neutrophils, $105$ macrophages, $386$ monocytes, $454$ granulocytes, and $833$ CD4 T cells – all collected from a single donor. In this controlled setting, where cell types were known, we succeeded in estimating the overdispersion of $8,\!333$ genes for each cell type using the variance stabilizing transformation implemented in the sctransform package (Choudhary and Satija,, 2022). Figure 5C illustrates the comparison of gene overdispersion when estimated solely in neutrophils versus the overdispersion estimated for the same genes in the other four cellular populations. Root Mean Squared Error (RMSE) values were computed to quantify the agreement between estimations. Our findings reveal that overdispersion is specific to each cell population, as evidenced by a notable deviation from the diagonal and relatively high RMSE values. This underscores the challenge of accurately estimating this parameter without prior knowledge of the true mixture underlying the data. Combined with our simulation studies, these results demonstrate how Type I error can easily be inflated in real-life applications of data thining for scRNA-seq data analysis, particularly due to the difficulty in providing an unbiased estimate of gene overdispersion.